It takes Mohammed 30 mins to get to work, and 50 minutes to return. He goes 6mph slower on the return. Find out how far she has to travel.
Answer by fractalier(6550)

Call his rate going to work, x.
Then his rate coming home is x - 6.
Expressing the times in hours, and using rt=d, we get
(1/2)x = (5/6)(x - 6)
Multiply by six to clear fractions...
3x = 5(x - 6)
3x = 5x - 30
-2x = -30
x = 15 mph
x-6 = 9mph
The distance to work is (1/2)(15) = 7.5 miles